JP Morgan Chase writes a forward contract to sell 10 million pounds sterling at the delivery price of $1.6247 per pound in 30 days’ time. At the end of the contract the spot exchange rate was $1.65 per pound. What was the profit to JP Morgan Chase?

The current spot price of a non-dividend paying stock is $65, while the simple interest rate is 4.50 percent per year. Consider a forward contract written on this stock with a maturity of 90 days. a) Calculate the forward price. b) What is the value of the contract? c) General Electric wants a 90-day forward contract with the delivery price set at $60. What is the value of this contract?
Answer #1

Given: Delivery price: 1 GBP= $1.627

spot exchange: 1 GBP = $1.65

Amount of currency to be sold: 10 million= 10,000,000 GBP

Profit = (difference between the rates) * Currency Amount

= (1.65-1.627) * 10,000,000 GBP

Profit = 230,000 GBP

Next:

a) Given : Spot price (S) = $65

Interest rate (r) = 4.5% per year

Maturity period (t) = 90 days = 90/360 year = 0.25 year

Using formula:

Forward Price F=S×(1+r)t

Putting values in the equation: F = 65* (1+ 0.045)(0.25)

F = 65.7192 Ans

b) Value of a Forward contract V =S − F /(1+r)

here S = spot price = $65 and F is the forward price we just calculated in the previous step.

putting values in the equation:

V = 65 - 65.7192/(1+0.045) = $2.1107 Ans

c) Again given t = 90 days =90/360 year =0.25 year

delivery price =$ 60

taking r = 0.045

using formula V =S − F /(1+r)

where S = $65 and F = $60

solving we get :

V = $7.5837